a. Use the chemical half reactions for free energy changes (deltaG0 standard conditions of pH=7.0 and 25o) to determine the free energy change from hydrogen (H2) oxidation by molecular oxygen (O2) to produce H2O.

b. Is the overall Redox reaction endothermic or exothermic?

c. Compare the value for the change in free energy in part (a) with the value for the change in free energy resulting from the oxidation of glucose by molecular oxygen that occurs in biological systems.

d. Can one tell from the results of a thermodynamic analysis if the rate of the glucose + O2 reaction is faster than the rate of the H2 + O2 reaction?
Answer #1

A)

2H2 (g) ---> 4 H+ (aq) + 4e- E0  = 0 Volts [occurs at cathode]

O2(g) + 4 H+(aq) + 4 e  arrow_right.gif 2 H2O [ccours at anone] E0  = -1.229 V

Ecell = Ecathode - Eanode

= (0) - (-1.229)

= 1.229 V

delta G = -nF E0cell [where n = number of electrons involved in the reaction, 1 faraday = (96458 C/mol e- ]

= - 4 * (96458 C/mol e-) * 1.229

= -474187.528 KJ

B). Since delta G value is negative value its exothermic reaction

C)

C6H12O6(s) + 6 O2(g) → 6 CO2(g) + 6 H2O(l)

The free energy for the oxidation of glucose to CO₂ and water is -2870.1448 kJl/mol (obtained from literature = -686 kcal/mol )

Comparing this with value obatined in part (a), the value for the change in free energy resulting from the oxidation of glucose by molecular oxygen is less than that of oxidation of hydrogen to yield water

part (d)

Based on the delta G values for both reactions, its evident that oxidation of hydrogen to yield water is more faster and spontaneous than the oxidation of glucose, since more negative is the delta G value more spontaneous the recation is.